An experiment was conducted to determine the effectiveness of a method designed to remove oil wastes found in soil. Three contaminated soil samples were treated. After 95 days, the percentage of contamination removed from each soil sample was measured, with a mean of 49.3 percent and a standard deviation of 1.5 percent. If we wished to narrow the boundary around μ for a 98 percent confidence interval to within 0.5 percent, how many soil samples should be in our experiment.

437

3

9

6

33
Answer #1

Correct option

n=[talpha/2*sd)/E]^2=(6.965*1.5/0.5)^2=436.6=437

Answer 437

Please upvote if this helps you out. Thank You!